A 100-W light bulb is connected to a 110-V source. What is the resistance of this bulb?

a) 121 O
b) 8.00 x 10-3 O
c) 6.00 x 10-3 O
d) 242 O
e) 100 O
Answer #1
P=V^2/R
R=V^2/P=121O
a) 121 O
answered by: Brandan
Answer #2

DATA
P=100W
E=110V

P=IE THEN I=P/E=100W/110V
I=0.909 A

I=E/R
R=E/I
R=110V/0.909A
R=121 Ω

answered by: Pierre Romo
Answer #3
a) 121 O
answered by: Jay111